OpenStudy (anonymous):

Jack usually mows his lawn in 5 hours. Marilyn can mow the same lawn in 4 hours. How much time would it take to mow the lawn together?

OpenStudy (anonymous):

jack takes 5 hours Marilyn - 4 hours jack does 1/5 of job in 1 hour Marilyn does 1/4 of job in 1 hour. to gether they do 1/4 +1/5 of the job in 1 hour. 1/4 +1/5 = 9/20 of the job in 1 hour together they will do the job in 20 /9 hours = 2.2 hours